# Fast X-ray Transient detection with AXIS: application to Magnetar Giant Flares

### v1

@@ 4 R-processes from MGFs
MGFs have also been suggested as potential sites for rapid neutron capture process (r-process) nucleosynthesis, responsible for forming many of the heaviest elements in the Universe (Cehula et al., 2024). The synchrotron radio afterglow following the prompt gamma-ray flare indicate an associated sub-relativistic baryon loaded ejecta component (e.g., Gelfand et al. 2005; Granot et al. 2006). Originating from the magnetar itself, the ejecta is sufficiently neutron-rich such that neutron captures proceed faster than beta-decays, enabling transmutation of light nuclei into heavier trans-iron species (Patel et al., 2025b). Recently Patel et al. (2025a) demonstrated direct evidence for this process, utilizing the archival data from the anticoincidence shield on the INTEGRAL satellite (Mereghetti et al., 2005) taken during the 2004 MGF from SGR 1806-20.
@@
Figure 6: X-ray predictions below at t = $10^{3}$ s, $10^{5}$ s, $10^{7}$ s (4 mo), and $10^{9}$ (32 yrs) for a source at 10 kpc. This is equivalent to Figure 3 of Patel et al. (2025a) but for different timescales after the event.
We can consider the single line emission. Firstly, we look at the case of 32 years after the event (bottom-right panel of Fig. 6) to see if AXIS, once launched could attempt an observation of the late r-process emission from SGR 1806-20, we focus on the $\sim 6.1$ keV line with a flux of $\sim 36$ ph/cm2/s, corresponding to an energy flux of $\sim 5.7\times 10^{-24}$ erg/cm2/s. Inserting this flux in webPIMMS we predict an AXIS photon rate of $3.56\times 10^{-13}$ counts/s with AXIS FOV-average in the 6–6.2 keV range. This number is $\sim 5$ orders of magnitude lower than the estimated background, suggesting that AXIS would not be able to detect such late-time emission.
@@ 5 Conclusions
Magnetar giant flares provide a rare but powerful probe of extreme physics, from extreme magnetic field dissipation and underlying trigger mechanisms, synthesis of heavy elements via r-process nucleosynthesis, and quantum electrodynamic processes (Burns et al., 2025). In this study, we assessed the detectability of MGFs with AXIS across multiple phases of their emission.
Serendipitous detections of short spikes are improbable within the AXIS field of view over the prime mission lifetime due to their intrinsic hardness and short duration, though bright events could in principle be observed out to several hundred Mpc. More broadly, we outlined a general workflow for estimating transient detectability with any instrument, combining sensitivity, volumetric rates, spectral assumptions, and mission efficiency.
@@ ACKNOWLEDGEMENTS
We wish to acknowledge the AXIS team for providing the instrument response functions and sensitivity estimates, crucial to carry out this study. The material is based upon work supported by NASA under award numbers 80GSFC21M0002 and 80GSFC24M0006. This work has made use of the NASA Astrophysics Data System.

### v2

@@ 4 R-processes from MGFs
MGFs have also been suggested as potential sites for rapid neutron capture process (r-process) nucleosynthesis, responsible for forming many of the heaviest elements in the Universe (Cehula et al., 2024). The synchrotron radio afterglow following the prompt gamma-ray flare indicates an associated sub-relativistic baryon-loaded ejecta component (e.g., Gelfand et al. 2005; Granot et al. 2006). Originating from the magnetar itself, the ejecta are sufficiently neutron-rich such that neutron captures proceed faster than beta-decays, enabling transmutation of light nuclei into heavier trans-iron species (Patel et al., 2025b). Recently Patel et al. (2025a) demonstrated direct evidence for this process, utilizing the archival data taken during the 2004 MGF from SGR 1806-20 with the anticoincidence shield on the INTEGRAL satellite (Mereghetti et al., 2005), the rear detectors on RHESSI (Boggs et al., 2007), and the Konus gamma-ray spectrometer on Wind (Frederiks et al., 2007).
@@
Figure 6: X-ray predictions below at t = $10^{3}$ s, $10^{5}$ s, $10^{7}$ s (4 mo), and $10^{9}$ (32 yrs) for a source at 10 kpc. Equivalent to Figure 3 of Patel et al. (2025a) but for different timescales after the event.
We can consider the single line emission. Firstly, we look at the case of 32 years after the event (bottom-right panel of Fig. 6) to see if AXIS, once launched could attempt an observation of the late r-process emission from SGR 1806-20, we focus on the $\sim 6.1$ keV line with a flux of $\sim 36\times 10^{-16}$ ph/cm2/s, corresponding to an energy flux of $\sim 5.7\times 10^{-24}$ erg/cm2/s. Inserting this flux into webPIMMS we predict an AXIS photon rate of $3.56\times 10^{-13}$ counts/s with AXIS FoV-average in the 6–6.2 keV range. This number is $\sim 5$ orders of magnitude lower than the estimated background, suggesting that AXIS would not be able to detect such late-time emission.
@@ 5 Conclusions
Magnetar giant flares provide a rare but powerful probe of extreme physics, including magnetic-field dissipation and triggering mechanisms, synthesis of heavy elements via r-process nucleosynthesis, and quantum electrodynamic processes (Burns et al., 2025). In this study, we assessed the detectability of MGFs with AXIS across multiple phases of their emission.
Serendipitous detections of short spikes are improbable within the AXIS field of view over the prime mission lifetime given their short duration and the assumption that the hard gamma-ray spectrum extends into the soft X-ray band, though bright events could in principle be observed out to several hundred Mpc, and the soft X-ray flux could be enhanced by a thermal component from an expanding fireball, as observed in short bursts and magnetar giant flare spikes. More broadly, we outlined a general workflow for estimating transient detectability with any instrument, combining sensitivity, volumetric rates, spectral assumptions, and mission efficiency.
@@ ACKNOWLEDGEMENTS
We wish to thank the anonymous referee whose feedback helped improve the manuscript. We also wish to acknowledge the AXIS team for providing the instrument response functions and sensitivity estimates, crucial to carry out this study. The material is based upon work supported by NASA under award numbers 80GSFC21M0002 and 80GSFC24M0006. This work has made use of the NASA Astrophysics Data System.